OpenStudy (anonymous):

The function f(x) = 10(5)x represents the growth of a lizard population every year in a remote desert. Crista wants to manipulate the formula to an equivalent form that calculates every half-year, not every year. Which function is correct for Crista's purposes?

OpenStudy (anonymous):

pleae help me!!!

OpenStudy (anonymous):

please**

OpenStudy (anonymous):

i really dont understand how to solve this

OpenStudy (jdoe0001):

as far as I can tell f(x) means, every "x" value is 1 year so f(2) is 2 years growth f(3) 3 years growth and so on to get half of that, wouldn't that be x/2? so \(\bf f\left( \frac{2}{2} \right)=f(1year)\qquad f\left( \frac{3}{2} \right)=f(1\frac{1}{2}year)\) and so on, thus "x" gets spliced in two, giving the two halves

OpenStudy (anonymous):

what does that mean?

OpenStudy (jdoe0001):

well.. do you know what the function is doing?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

those are the answers i am given and it all looks like gibberish

OpenStudy (jdoe0001):

I assume you've covered what a function is already? and what an input and output is?

OpenStudy (anonymous):

yes? but I am not clear on it. I am English math is scary to me

OpenStudy (jdoe0001):

on sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

by the time you read this i will have gone ill be back but my parents are bugging me. so i must leave.

OpenStudy (anonymous):

oh nvm

OpenStudy (jdoe0001):

\(\large { \begin{array}{llll} \textit{lizard population}&f(x)=10(5)^x \\\hline\\ year&amount \\\hline\\ 1&f(1)\to 10(5)^1\\ 2&f(2)\to 10(5)^2\\ 3&f(3)\to 10(5)^3\\ 4&f(4)\to 10(5)^4\\ 5&f(5)\to 10(5)^5\\ ...&... \end{array} }\) see how the population is increasing as the years go by?

OpenStudy (anonymous):

yes the year is equal to the power

OpenStudy (jdoe0001):

the "x" accounts for the year amount so, if you want 6months, or half a year split the "x" or the year, in half, or x/2 now if you make f(x) look like \(\bf f\left( \frac{x}{2} \right)\) what would the expression on the right-hand-side look like then? \(\bf 10(5)^x\) will look like ?

OpenStudy (anonymous):

10/2 ^(5) ^x?

OpenStudy (jdoe0001):

hmm? notice you're changing the "years" the 10 is not the years though

OpenStudy (jdoe0001):

\(\large f(year)=10(5)^\textit{year}\qquad f\left( \frac{year}{2} \right)=?\)